OpenStudy (anonymous):

The school band is scheduled to march in the annual parade, but the band’s starting point is changed the day before the parade. The band director calls three band members. Each band member calls three other band members. Then these band members each call three members. How many band members, including the band director, are notified of the new starting point?

OpenStudy (mayaal):

I guess its 162

OpenStudy (anonymous):

3+3+3=9*3=27+1 So 28 are notified

OpenStudy (mayaal):

1 band director,3 band member,they call 9 other band members,then they call 27 others.Now add them together...

OpenStudy (mayaal):

So 40 are notified...

OpenStudy (anonymous):

A. 12 members B. 13 members C. 20 members D. 40 members

OpenStudy (mayaal):

Look its D.40

OpenStudy (anonymous):

:P Oops.

OpenStudy (mayaal):

there is no choice for 28

OpenStudy (mayaal):

In which grade r u @BraydonG ?

OpenStudy (anonymous):

8th Algebra I miss calculated in one part but I fixed it

OpenStudy (mayaal):

i am in 8th grade too but in geometry well its ok you will understand it when you will study geometry....Good job u were near..... :)